OpenStudy (anonymous):

If a stone is tossed from the top of a 210 meter building, the height of the stone as a function of time is given by h(t) = -9.8t2 – 10t + 210, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth’s place; include units in your answer.

Nnesha (nnesha):

when stone hits the ground height would be zero  so replace h(t) by 0 and then use quadratic formula to find t \[\huge\rm \frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\]

Nnesha (nnesha):

\[\huge\rm \color{Red}{0} = -9.8t2 – 10t + 210 ~~~~\color{red}{h(t) =0}\]